Shutdown stsetch SNF 2010-06-22 11:20:07: Cooling He leak.

eenriquez at snf.stanford.edu eenriquez at snf.stanford.edu
Tue Jun 22 14:18:42 PDT 2010


Replaced the lip seal o-ring.  Now the helium flow is 8.5 - 9.5 with the setpoint of 9.7.  Its better but its not repaired yet.  Will take it off shutdown just in case somebody wants to run at the current helium flow level.




More information about the stsetch-pcs mailing list